# Deferred Commitment Decoding for Diffusion Language Models with Confidence-Aware Sliding Windows

### v1

@@ 5.3 DCD’s Combination with KV Cache
To accelerate DLM inference, we integrate prefix and dual caching into the DCD algorithm, following Fast-dLLM wu2025fast. Inspired by dKV-Cache-Greedy ma2025dkv, the active interval without caching is slightly extended beyond the decoded tokens from the current and previous steps. Formally, it is defined as:
@@
We then define the prefix of the generation slot as the tokens preceding $\mathcal{W}^{(t)}$, and the suffix as the tokens following $\mathcal{W}^{(t)}$. The prefix cache temporarily stores the prefix, while the dual cache stores both the prefix and suffix. Additionally, to ensure a fair comparison with block-based cache refreshing, we rebuild the cache after $B^{\prime}$ masked tokens have been decoded since the previous cache refresh.
@@ 6 Experiments
Table 1: Experimental results. For each experiment, we report its overall metrics (pass@1, accuracy, etc.). We also report the total seconds for running 5 benchmarks within one line. The best result of certain model and task is bolded and second-best is underlined.
@@
Fast-dLLM -v2-7B Avg. Metric +0.62 Avg. Time -8.3% None Block-based 9993 56.7 48.4 50.8 74.5 62.5
None Sub-block-based 11228 61.0 50.2 54.6 77.6 62.8
@@
Figure 2: Confidence distributions of LLaDA-8B-Instruct, Dream-v0-Base-7B, and Fast-dLLM-v2-7B on GSM8K with DCD and (sub-)block-based decoding using dual cache. We use a log10 scale for clarity. The DCD algorithm yields fewer low-confidence decoding steps across these models.
Figure 3: Ablation studies on LLaDA-8B-Instruct on the MATH500 task with dual-cache DCD. We vary $s_{\text{max}}$, $s_{\text{init}}$, and $\tau_{\text{conf}}$, and evaluate task accuracy and decoding time.
@@ 6.1 Experimental Setup
To validate the effectiveness of the DCD algorithm, we evaluate four pretrained DLMs: LLaDA-8B-Instruct nie2025large, Dream-v0-Instruct-7B, Dream-v0-Base-7B ye2025dream, and Fast-dLLM-v2-7B wu2025fast2. The first three models adopt full attention, while the last model uses semi-causal attention.
@@
For each model, we evaluate coding benchmarks including HumanEval humaneval and MBPP mbpp, mathematical reasoning benchmarks including MATH500 math500 and GSM8K gsm8k, and the instruction-following benchmark IFEval ifeval. The Dream-v0-Base-7B model is excluded from IFEval because it is not instruction-aligned. We do not evaluate multiple-choice QA benchmarks mmlu; gpqa, as they primarily measure token-level log-probabilities rather than decoding quality.
@@
For full-attention DLMs, the parallel block-based decoding of Fast-dLLM wu2025fast serves as the baseline under three cache configurations: no cache, prefix cache, and dual cache. For the semi-causal DLM Fast-dLLM-v2-7B wu2025fast2, sub-block-based decoding with no cache and with dual cache within each large block is used as the primary baseline, while vanilla block-based decoding without sub-block structures is used as an additional baseline. For each model, benchmark, and cache configuration, DCD is compared against these baselines as well as other training-free DLM decoding strategies such as dKV-Cache ma2025dkv and AdaBlock-dLLM lu2025adablock, when available.
@@
For full-attention DLMs, we set the generation slot length to $L=512$, window parameters to $s_{\text{init}}=16$ and $s_{\text{max}}=128$, cache parameters to $B^{\prime}=32$ and $r=2$, and the confidence threshold to $\tau_{\text{conf}}=0.9$. For the semi-causal DLM, we set $s_{\text{init}}=8$ and $s_{\text{max}}=\infty$, while keeping all other parameters unchanged. All baselines use a block size of $B=32$ across all four models, with a sub-block size of $b=8$ for Fast-dLLM-v2-7B wu2025fast2. The decoding temperature is set to 0 and the batch size is set to 1 for all experiments. Additional details are provided in Appendix C.
@@ 6.2 Main Results Analysis
Table 1 reports the main experimental results across multiple diffusion language models, tasks, and KV-cache configurations. Overall, Deferred Commitment Decoding (DCD) consistently outperforms block-based and sub-block-based decoding across most settings, demonstrating strong robustness across tasks and model architectures. On average, DCD improves evaluation metrics by +1.39% while reducing decoding time by 4.4% compared to block-based (or sub-block-based) baselines across all models. The average improvement for each model against baseline is highlighted in colored text in Table 1.
These gains are observed across diverse tasks, including mathematical reasoning, code generation, and instruction following, and across both full-attention and semi-causal DLMs. Among all models, Dream-v0-Instruct-7B benefits the most from DCD, achieving the largest average improvement of 2.53 points. Running the MBPP benchmark with Dream-v0-Instruct-7B and dual cache yields the most significant improvement, with a 9.0% increase in pass@1. In contrast, Fast-dLLM-v2-7B shows the smallest improvement of 0.62 points. This outcome is expected, as DCD currently operates only at the sub-block level for semi-causal models and does not modify the macro-level block-based decoding paradigm imposed by their attention structure. Nevertheless, DCD still consistently outperforms sub-block-based decoding, which aligns with our theoretical analysis.
Compared with AdaBlock lu2025adablock, DCD achieves better performance (average metric improvements of +0.28% for LLaDA-8B-Instruct and +1.20% for Dream-v0-Base-7B) and substantial speedups (average decoding time reductions of 19% and 71%, respectively), demonstrating the effectiveness of the deferred commitment mechanism. Compared with dKV-Cache-Greedy ma2025dkv, DCD substantially outperforms it in terms of accuracy, despite employing a similar KV-cache strategy.
We note that in a small number of cases, DCD performs slightly worse than block-based decoding. We attribute these regressions to the inherent stochasticity of training-free decoding and the intrinsic difficulty of certain tokens, which may remain ambiguous even with extended context.
@@ 6.3 Evidence of BICT Mitigation
Across all models, DCD substantially reduces the frequency of extremely low-confidence decoding steps compared to block-based or sub-block-based decoding. Such low-confidence events directly reflect the BICT phenomenon that DCD is designed to address. This reduction provides a clear explanation for the observed accuracy improvements, particularly on reasoning-intensive tasks.
@@ 6.4 Ablation Studies
Figure 3 presents ablation studies on LLaDA-8B-Instruct evaluated on MATH500 with dual cache enabled, analyzing the impact of three key hyperparameters in Deferred Commitment Decoding (DCD): the maximum window size $s_{\max}$, the initial window size $s_{\text{init}}$, and the confidence threshold $\tau_{\text{conf}}$.
@@
Effect of $\tau_{\text{conf}}$.
The confidence threshold regulates how aggressively tokens are deferred. Accuracy improves as $\tau_{\text{conf}}$ increases from lower values, but degrades when $\tau_{\text{conf}}=1$ (i.e., top-1 confidence decoding), as this setting becomes fully deterministic and loses flexibility. In contrast to window parameters, $\tau_{\text{conf}}$ exhibits a clearer positive correlation with decoding time.
Overall, accuracy exhibits a clear unimodal trend as $s_{\max}$, $s_{\text{init}}$, and $\tau_{\text{conf}}$ increase in this setting. Based on these ablation results, we select appropriate hyperparameters and apply them consistently across all experiments.
@@ 7 Conclusion
In this work, we investigate a fundamental limitation of block-based diffusion decoding for language models, which we formalize as Boundary-Induced Context Truncation. We identify suboptimal token commitment, whereby tokens that would otherwise benefit from nearby future context are forced to commit at block boundaries, leading to low-confidence predictions and degraded generation quality. To address this issue, we propose Deferred Commitment Decoding, a simple, training-free decoding strategy that replaces fixed block boundaries with a confidence-aware sliding window. By deferring uncertain tokens until sufficient context becomes available, DCD enables more effective utilization of local bidirectional context without sacrificing KV-cache compatibility. Extensive experiments across multiple DLMs, benchmarks, and caching configurations demonstrate that DCD consistently improves generation accuracy by 1.39% with comparable time on average compared to fixed block-based baselines, with the maximum improvement reaching 9.0%. These results demonstrate the superiority of DCD for DLM inference.
Limitations
A major limitation of the proposed DCD method arises in semi-causal DLMs, where we only modify the sub-block structure and consequently observe smaller improvements than in full-attention DLMs. We hope that future work will adapt semi-causal DLM architectures to better accommodate the DCD mechanism and achieve stronger performance across a wider range of benchmarks.
@@ Appendix A Experimental Environment
For comparison experiments between block-based decoding and DCD, we allocate one NVIDIA A100 80G GPU and eight Intel(R) Xeon(R) Platinum 8358 CPUs for each experiment. For AdaBlock experiments, we allocate one NVIDIA A800 GPU and sixteen Intel(R) Xeon(R) Platinum 8378A CPUs for each experiment.
@@
We use Ubuntu Linux, CUDA 12.0, Python 3.10, and lm-eval-harness 0.4.8 for all experiments.
Appendix B A Case Study of the BICT Phenomenon
Figure 4: MBPP Task #129 solved by block-based decoding and the DCD algorithm using Dream-v0-Instruct-7B.
Figure 4 illustrates the BICT phenomenon in block-based decoding. This example corresponds to the 129th test case of the MBPP benchmark, generated by Dream-v0-Instruct-7B without any caching. The prompt for this task is: “Write a function to extract elements that occur singly in the given tuple list.” The test cases are:
@@ Appendix B A Case Study of the BICT Phenomenon
The error occurs at decoding step 21, as illustrated in the left part of Figure 4. The $\langle\text{MASK}\rangle$ at a critical position (the rightmost token of line 5) is adjacent to the first block boundary and suffers from truncated context. As a result, it is incorrectly decoded as “tup” with low confidence.
@@ Appendix C Details about Main Experiments
All evaluation scores are generated using lm-eval-harness 0.4.8 and the code-cleaning suite in the Fast-dLLM codebase. Specifically:
@@
Appendix D Time Consumption of Each Experiment
To better understand the DCD method, we record the time consumption for each benchmark experiment. According to Table 2, DCD completes benchmarks in comparable and even slightly less time than the baseline method, demonstrating its efficiency against traditional methods.
@@ Appendix D Time Consumption of Each Experiment
Fast-dLLM -v2-7B None Block-based 561 1487 2168 2831 2946
None Sub-block-based 615 1566 2650 3285 3112

### v2

@@ 5.3 DCD’s Combination with KV Cache
To accelerate DLM inference, we integrate prefix and dual caching into the DCD algorithm, following Fast-dLLM Wu et al. (2025b). Inspired by dKV-Cache-Greedy Ma et al. (2025), the active interval without caching is slightly extended beyond the decoded tokens from the current and previous steps. Formally, it is defined as:
@@
We then define the prefix of the generation slot as the tokens preceding $\mathcal{W}^{(t)}$, and the suffix as the tokens following $\mathcal{W}^{(t)}$. The prefix cache temporarily stores the prefix, while the dual cache stores both the prefix and suffix. Additionally, to ensure a fair comparison with block-based cache refreshing, we rebuild the cache after $B^{\prime}$ masked tokens have been decoded since the last cache refresh.
@@ 6 Experiments
Table 1: Experimental results. For each experiment, we report its overall metrics (pass@1, accuracy, etc.). We also report the total seconds for running 5 benchmarks within one line. The best result of certain model and task is bolded and second-best is underlined. The (sub-)block-based decoding serves as DCD’s baseline and other algorithms serve as comparison works.
@@
Fast-dLLM -v2-7B Avg. Metric +0.62 Avg. Time -8.3% None Sub-block-based 11228 61.0 50.2 54.6 77.6 62.8
None DCD 10116 62.8 48.6 53.4 77.9 64.0
@@
None Block-based 9993 56.7 48.4 50.8 74.5 62.5
NBDiff Avg. Metric +5.22 None Sub-block-based - 82.3 78.2 80.1 87.4 40.1
None DCD 526163 81.7 80.2 84.4 91.3 56.6
Figure 4: Ablation studies. We vary $s_{\text{max}}$, $s_{\text{init}}$, and $\tau$, and evaluate task accuracy and decoding time on LLaDA-8B-Instruct on the MATH500 task with dual-cache DCD.
@@ 6.1 Experimental Setup
To validate the effectiveness of the DCD algorithm, we evaluate full-attention pretrained DLMs including LLaDA-8B-Instruct Nie et al. (2025), Dream-v0-Instruct-7B, and Dream-v0-Base-7B Ye et al. (2025); and semi-causal DLMs including Fast-dLLM-v2-7B Wu et al. (2025a) and NBDiff Tian et al. (2025).
@@
For each model, we evaluate coding benchmarks including HumanEval Chen et al. (2021) and MBPP Austin et al. (2021b), mathematical reasoning benchmarks including MATH500 Lightman et al. (2023) and GSM8K Cobbe et al. (2021), and the instruction-following benchmark IFEval Zhou et al. (2023). The Dream-v0-Base-7B model is excluded from IFEval because it is not instruction-aligned. We do not evaluate multiple-choice QA benchmarks Hendrycks et al. (2020); Rein et al. (2024), as they primarily measure token-level log-probabilities rather than decoding quality.
@@
For full-attention DLMs, the parallel block-based decoding of Fast-dLLM Wu et al. (2025b) serves as the baseline under three cache configurations: no cache, prefix cache, and dual cache. For the semi-causal DLM Fast-dLLM-v2-7B Wu et al. (2025a), sub-block-based decoding with no cache and with dual cache within each large block is used as the primary baseline, while vanilla block-based decoding without sub-block structures is used as an additional baseline. For each model, benchmark, and cache configuration, DCD is compared against these baselines as well as other training-free DLM decoding strategies such as dKV-Cache Ma et al. (2025), AdaBlock-dLLM Lu et al. (2025), CCD Chen et al. (2025), and Prophet Li et al. (2025) when available.
@@
All experiments use block size $B=32$, sub-block size $b=8$ and $\text{batchsize}=1$. To align with the configurations of baseline works, decoding certainty is measured by negative entropy for NBDiff Tian et al. (2025) and by confidence for the remaining DLMs Wu et al. (2025b, a), with a unified threshold $\tau=0.9$. For full-attention DLMs, we set $L=512$, $s_{\text{init}}=16$, $s_{\text{max}}=128$, $B^{\prime}=32$ and $r=2$. For the semi-causal DLM, we set $s_{\text{init}}=8$ while keeping all other parameters unchanged. The DBE is not enabled in main results and discussed in Section 6.5 with $e_{\text{step}}=4,e_{\text{max}}=16$ and $\tau_{\text{low}}=0.4$ for confidence and $\tau_{\text{low}}=-0.5$ for negative entropy.
@@ 6.2 Main Results Analysis
Table 1 reports the main experimental results across multiple diffusion language models, tasks, and KV-cache configurations. Overall, Deferred Commitment Decoding (DCD) consistently outperforms block-based and sub-block-based decoding across most settings, demonstrating strong robustness across tasks and model architectures. On average, DCD improves evaluation metrics by +1.73% while reducing decoding time by 4.4% compared to block-based (or sub-block-based) baselines with regard to the same model, task and cache configuration. The average improvement for each model over the baseline is highlighted in colored text in Table 1.
DCD’s gains over (sub-)block-based baselines are observed across diverse tasks—including mathematical reasoning, code generation, and instruction following—and across both full-attention and semi-causal DLMs. Among all models, LLaDA-8B-Instruct, Dream-v0-Instruct-7B, and Dream-v0-Base-7B exhibit significant improvements in accuracy and moderate speedups in time consumption, showing that our DCD algorithm works effectively at the whole-sequence level. NBDiff benefits the most from DCD, achieving the largest average improvement of 5.22%, as it is trained with multiple block sizes and thus aligns well with sliding windows. Running the IFEval benchmark with NBDiff yields the most significant improvement, with a 16.5% increase in accuracy. In contrast, Fast-dLLM-v2-7B shows the smallest improvement of 0.62% because it is trained with limited flexibility with variable decoding ranges. Nevertheless, DCD still outperforms sub-block-based decoding, which aligns with our theoretical analysis.
Compared with AdaBlock Lu et al. (2025), DCD achieves better performance (average metric improvements of +0.28% for LLaDA-8B-Instruct and +1.20% for Dream-v0-Base-7B) and substantial speedups (average decoding time reductions of 19% and 71%, respectively), demonstrating the effectiveness of the deferred commitment mechanism. Compared with dKV-Cache-Greedy Ma et al. (2025), DCD substantially outperforms it in terms of accuracy, despite employing a similar KV-cache strategy. Compared with CCD Chen et al. (2025), which uses a custom cache to store historical data, the best results of DCD surpass it by 1.90% on average. More importantly, DCD outperforms Prophet Li et al. (2025) by a large margin because the latter algorithm contradicts DCD by performing early decoding, which substantially degrades DLMs’ inference performance.
We note that in a small number of cases, DCD performs slightly worse than (sub-)block-based decoding and other methods. We attribute these regressions to the inherent stochasticity of training-free decoding and the intrinsic difficulty of certain tokens, which may remain ambiguous even with extended context.
@@ 6.3 Evidence of BICT Mitigation
Across all models, DCD substantially reduces the frequency of extremely low-certainty decoding steps compared to block-based or sub-block-based decoding. Such low-certainty events directly reflect the BICT phenomenon that DCD is designed to address. This reduction provides a clear explanation for the observed accuracy improvements, particularly on reasoning-intensive tasks.
@@ 6.4 Ablation Studies
Figure 4 presents ablation studies on LLaDA-8B-Instruct evaluated on MATH500 with dual cache enabled, analyzing the impact of three key hyperparameters in DCD: the maximum window size $s_{\max}$, the initial window size $s_{\text{init}}$, and the certainty threshold $\tau$.
@@
Effect of $\tau$.
The certainty threshold regulates how aggressively tokens are deferred. Accuracy improves as $\tau$ increases from lower values but degrades when $\tau=1$ (i.e., top-1 confidence decoding), as this setting becomes fully deterministic and loses flexibility. In contrast to window parameters, $\tau$ exhibits a clearer positive correlation with decoding time.
Overall, accuracy exhibits a clear unimodal trend as $s_{\max}$, $s_{\text{init}}$, and $\tau$ increase in this setting. Based on these ablation results, we select appropriate hyperparameters and apply them consistently across all experiments.
6.5 The Effects of Dynamic Block Extention
Table 2 reports the performance of DBE-patched DCD across five benchmarks for semi-causal DLMs. Despite minimal computational overhead, NBDiff consistently benefits from DBE across various tasks, achieving an average improvement of 1.14%, as expected. This is because NBDiff’s training data includes generation slots of varying block sizes. In contrast, Fast-dLLM-v2 is trained with a fixed block size of $B=32$, and thus fails to benefit from DBE, instead suffering a 1.30% drop in metrics. Overall, we suggest that semi-causal DLMs should be trained with variable block sizes to mitigate the BICT problem and enable the DBE mechanism.
Table 2: Experimental results for DBE-patched DCD. The comparison baselines are vanilla DCD without cache for two models.
Item NBDiff Fast-dLLM-v2
HumanEval $83.5^{{\color[rgb]{1,0,0}+1.8}}$ $58.5^{{\color[rgb]{0,0.49609375,0}-4.3}}$
MBPP $80.9^{{\color[rgb]{1,0,0}+0.7}}$ $47.2^{{\color[rgb]{0,0.49609375,0}-1.4}}$
MATH500 $86.2^{{\color[rgb]{1,0,0}+1.8}}$ $54.0^{{\color[rgb]{1,0,0}+0.6}}$
GSM8K $90.9^{{\color[rgb]{0,0.49609375,0}-0.4}}$ $78.4^{{\color[rgb]{1,0,0}+0.5}}$
IFEval $58.4^{{\color[rgb]{1,0,0}+1.8}}$ $62.1^{{\color[rgb]{0,0.49609375,0}-1.9}}$
Avg. Metric $79.98^{{\color[rgb]{1,0,0}+1.14}}$ $60.04^{{\color[rgb]{0,0.49609375,0}-1.30}}$
Time $604526^{{\color[rgb]{1,0,0}+14.9\%}}$ $10757^{{\color[rgb]{1,0,0}+6.3\%}}$
@@ 7 Conclusion
In this work, we investigate a fundamental limitation of block-based diffusion decoding for language models, which we formalize as Boundary-Induced Context Truncation. We identify suboptimal token commitment—whereby tokens that would otherwise benefit from nearby future context are forced to commit at block boundaries—leading to low-certainty predictions and degraded generation quality. To address this issue, we propose Deferred Commitment Decoding, a simple, training-free decoding strategy that replaces fixed block boundaries with sliding windows. By deferring uncertain tokens until sufficient context becomes available, DCD enables more effective utilization of local bidirectional context without sacrificing KV-cache compatibility. Extensive experiments across multiple DLMs, benchmarks, and caching configurations demonstrate that DCD consistently improves generation accuracy by 1.73% on average with comparable decoding time relative to fixed block-based baselines, with the maximum improvement reaching 16.5%. These results demonstrate the superiority of DCD for DLM inference.
@@ Appendix A Experimental Environment
For comparison experiments between (sub-)block-based decoding and DCD on LLaDA-8B-Instruct, Dream-v0-Base-7B, and Fast-dLLM-v2-7B, we allocate one NVIDIA A100 80GB GPU and eight Intel(R) Xeon(R) Platinum 8358 CPUs for each experiment. For AdaBlock and NBDiff experiments, we allocate one NVIDIA A800 GPU and sixteen Intel(R) Xeon(R) Platinum 8378A CPUs for each experiment.
@@
We use Ubuntu Linux, CUDA 12, and Python 3.10 for all experiments. For LLaDA-8B-Instruct, Dream-v0-Base-7B, and Fast-dLLM-v2-7B, we follow their original papers and use lm-eval-harness as the evaluation suite. For NBDiff, we use opencompass as the evaluation suite.
Appendix B More Details about BICT Case Study
Figure 2 in main paper illustrates the BICT phenomenon in block-based decoding. This example corresponds to the 129th test case of the MBPP benchmark, generated by Dream-v0-Instruct-7B without any caching. The prompt for this task is: “Write a function to extract elements that occur singly in the given tuple list.” The test cases are:
@@ Appendix B More Details about BICT Case Study
The error occurs at decoding step 21, as illustrated in the left part of Figure 2. The $\langle\text{MASK}\rangle$ at a critical position (the rightmost token of line 5) is adjacent to the first block boundary and suffers from truncated context. As a result, it is incorrectly decoded as ”tup” with low confidence.
@@ Appendix C Details about Main Experiments
C.1 lm_eval_harness Part
For LLaDA-8B-Instruct, Dream-v0-Base-7B and Fast-dLLM-v2-7B, we use lm-eval-harness 0.4.8 and the code-cleaning suite in the Fast-dLLM codebase. Specifically:
@@
C.2 OpenCompass Part
For NBDiff, we use zero-shot evaluation with simple or no prompt engineering, as detailed below. The scoring logic is entirely based on https://github.com/open-compass/opencompass. These configurations are identical to those in the original paper.
  • HumanEval:
    Complete the following python code:
    {{text}}
  • MBPP:
    You are an expert Python programmer, and here is your task: {{text}}
    Your code should pass these tests:
    {{test_list[0]}}
    {{test_list[1]}}
    {{test_list[2]}}
    You should submit your final solution in the following format: ‘‘‘python
  • MATH500:
    {{text}}
[…]
  • GSM8K:
    Question: {{text}}
[…]
  • IFEval: No additional prompt.
Appendix D More Statistics of Each Experiment
D.1 Time Consumption
To better understand the DCD method, we record the time consumption for each benchmark experiment. According to Table 3, DCD completes benchmarks in comparable—and even slightly less—time than the baseline method, demonstrating its efficiency relative to traditional approaches. For dKV-Cache-Greedy, CCD, and Prophet, we copy their statistics from the original papers, as their time consumption data are unavailable.
@@ D.1 Time Consumption
Fast-dLLM -v2-7B None Sub-block-based 615 1566 2650 3285 3112
None DCD 561 1519 2253 2793 2990
@@
None Block-based 561 1487 2168 2831 2946
NBDiff None DCD 35215 81001 179530 135040 95467
D.2 Low-Certainty Decoding
To better understand the BICT phenomenon, we collect the number of low-certainty ($\text{confidence}<0.3$) decoding steps in each experiment. The results show that the DCD algorithm mitigates the BICT phenomenon by significantly reducing the number of low-certainty decoding steps, thereby improving performance.
Table 4: Low-certainty decoding steps of each experiment.
Model Cache Decoding Humaneval MBPP MATH500 GSM8K IFEval
LLaDA-8B -Instruct None Block-based 57 214 366 161 7450
None DCD 52 177 284 129 6848
Prefix Block-based 67 220 370 184 7870
Prefix DCD 50 269 330 137 6975
Dual Block-based 88 333 538 242 $81$
Dual DCD 82 245 421 155 7245
Dream-v0- Instruct-7B None Block-based 181 43 473 588 7481
None DCD 165 45 395 579 7381
Prefix Block-based 193 66 508 647 7935
Prefix DCD 159 54 446 640 7493
Dual Block-based 227 75 689 746 8304
Dual DCD 196 59 565 702 7914
Dream-v0- base-7B None Block-based 1393 951 1213 9111 -
None DCD 353 631 888 601 -
Prefix Block-based $14$ 1129 1408 8726 -
Prefix DCD 399 878 899 683 -
Dual Block-based 1499 1277 1477 6084 -
Dual DCD 471 971 894 791 -
Fast-dLLM -v2-7B None Sub-block-based 244 701 292 805 5679
None DCD 184 604 158 160 4512
Dual Sub-block-based 332 994 401 1119 6748
Dual DCD 174 690 175 181 4751
None Block-based 192 602 245 583 4861
D.3 Other Efficiency Metrics
The wall-clock time reported earlier is highly dependent on hardware-specific environments and exhibits limited generalization. Therefore, we report two additional metrics to assess the algorithmic efficiency of each experiment:
  • Average decoding steps. This metric measures the average number of decoding steps required to complete generation for prompts. It corresponds directly to the average number of forward propagations and constitutes the primary source of time consumption in DLM inference.
  • Average forward length. This metric measures the average number of tokens directly fed into the DLM for forward propagation. In some scenarios, certain tokens are cached and thus excluded from this count. Consequently, this metric is positively correlated with both time consumption and memory usage per decoding step.
As shown in Table 5 and Table 6, the average decoding steps of DCD are slightly lower (-5.0% on average) than those of the (sub-)block-based baselines, indicating a modest efficiency gain from its decoding strategy. Meanwhile, the average forward length under DCD is comparable (+6.0% on average) to that of the baselines across all models and tasks. Together, these results suggest that DCD incurs no additional computational overhead—in fact, its total inference time is on par with or slightly less than that of the (sub-)block-based approaches, consistent with the wall-clock time measurements.
Table 5: Average decoding steps of each experiment.
Model Cache Decoding Humaneval MBPP MATH500 GSM8K IFEval
LLaDA-8B -Instruct None Block-based 161.4 108.1 154.4 86.0 197.1
None DCD 160.6 108.1 149.2 85.0 201.2
Prefix Block-based 163.4 112.5 157.3 89.8 197.4
Prefix DCD 165.2 113.4 151.5 87.0 199.5
Dual Block-based 174.9 118.5 168.4 95.0 195.9
Dual DCD 169.5 113.3 162.1 90.4 196.8
Dream-v0- Instruct-7B None Block-based 88.5 28.0 150.1 55.7 176.7
None DCD 85.9 26.4 147.9 55.0 179.7
Prefix Block-based 88.1 28.8 152.0 56.5 172.5
Prefix DCD 85.4 26.6 149.5 55.6 169.2
Dual Block-based 86.0 30.0 161.2 57.6 164.1
Dual DCD 83.1 27.3 155.8 55.9 165.3
Dream-v0- base-7B None Block-based 129.7 120.2 169.9 64.9 -
None DCD 113.2 123.5 162.0 51.3 -
Prefix Block-based 132.8 130.1 160.8 65.8 -
Prefix DCD 119.6 136.4 153.4 52.2 -
Dual Block-based 132.5 138.3 163.2 68.1 -
Dual DCD 121.2 141.0 149.6 53.6 -
Fast-dLLM -v2-7B None Sub-block-based 157.0 130.1 216.4 108.3 253.4
None DCD 143.9 124.1 196.7 92.5 243.9
Dual Sub-block-based 157.9 136.9 224.5 112.7 258.2
Dual DCD 143.2 121.8 198.0 94.5 246.4
None Block-based 143.9 124.1 196.7 92.5 243.9
NBDiff None DCD 4787.1 6185.1 7166.3 2573.3 3376.2
Table 6: Average forward length of each experiment.
Model Cache Decoding Humaneval MBPP MATH500 GSM8K IFEval
LLaDA-8B -Instruct None Block-based 672.3 1259.9 674.6 1540.3 571.1
None DCD 672.5 1261.0 673.7 1540.9 571.1
Prefix Block-based 317.4 416.5 347.0 504.3 354.9
Prefix DCD 312.9 410.4 345.1 506.9 366.2
Dual Block-based 86.1 134.7 83.0 172.2 56.9
Dual DCD 88.8 139.0 81.5 180.8 72.8
Dream-v0- Instruct-7B None Block-based 681.4 1234.7 667.9 1532.5 576.9
None DCD 685.0 1241.2 668.1 1531.1 576.6
Prefix Block-based 435.3 550.1 375.2 552.8 368.6
Prefix DCD 436.4 572.6 373.3 560.2 383.3
Dual Block-based 69.1 116.8 78.1 148.0 57.5
Dual DCD 70.0 152.0 78.1 163.8 69.0
Dream-v0- base-7B None Block-based 646.0 1203.3 646.9 1511.5 -
None DCD 646.6 1203.1 647.7 1513.3 -
Prefix Block-based 405.2 472.5 344.3 526.4 -
Prefix DCD 415.4 470.3 344.2 558.9 -
Dual Block-based 65.7 167.5 90.5 121.6 -
Dual DCD 71.2 170.6 90.2 152.6 -
Fast-dLLM -v2-7B None Sub-block-based 32.9 37.3 32.6 41.2 32.2
None DCD 33.0 37.6 32.7 42.8 32.2
Dual Sub-block-based 15.5 19.3 18.0 25.7 13.5
Dual DCD 20.4 24.4 22.5 32.4 19.4
None Block-based 33.0 37.6 32.7 42.8 32.2
NBDiff None DCD 32.0 32.0 32.0 32.0 32.0